If x^64 = 64 then x^32=
(A)8 or -8
(B)12 or -12
(C)16
(D)32 or -32
(E)48

Answer by stanbon(48535)

If x^64 = 64 then x^32=
(A)8 or -8
(B)12 or -12
(C)16
(D)32 or -32
(E)48
-------------------------
x^64-64 = 0
Factor:
(x^32-8)(x^32+8) = 0
x^32 = 8 or x^32 = -8
==========================
Cheers,
Stan H.